OpenStudy (anonymous):

-3(5y-8)-y=-2(y-5) Help!!!

OpenStudy (nurali):

-3(5y-8)-y=-2(y-5) -15y+24-y=-2y+10 -16y+24=-2y+10 -16y+2y=10-24 -14y=-14 y=1

OpenStudy (anonymous):

Okay, so you distributed, then combined like terms, correct?

OpenStudy (anonymous):

Oh, okay I see. Thanks! That helped me out a lot!

OpenStudy (nurali):

Anytime.